# Self-Ensemble: Mitigating Confidence Distortion for Large Language Models

### v1

@@ Abstract
Although Large Language Models (LLMs) perform well in general fields, they exhibit a confidence distortion problem on multi-choice question-answering (MCQA), particularly as the number of answer choices increases. Specifically, on MCQA with many choices, LLMs suffer from under-confidence in correct predictions and over-confidence in incorrect ones, leading to a substantially degraded performance. To solve this problem, we propose Self-Ensemble in this work. Our method splits the choices into several groups and ensembles LLM predictions across these groups to reach a final decision. The advantage of Self-Ensemble is its plug-and-play nature, where it can be integrated into existing LLM architecture based on a designed attention mask and positional encoding, without requiring labeled datasets for parameter tuning. Experimental results on three LLMs and datasets demonstrate that Self-Ensemble comprehensively addresses the confidence distortion problem of LLMs, outperforming standard inference as well as baseline methods.
Self-Ensemble: Mitigating Confidence Distortion for
@@
Zicheng Xu1,3thanks: Equal contribution, ordered by rolling dices., Guanchu Wang211footnotemark: 1, Guangyao Zheng3, Yu-Neng Chuang1, Alexander Szalay3, Xia Hu1, Vladimir Braverman3 1Rice University, 2University of North Carolina at Charlotte, 3Johns Hopkins University Correspondence: vova@cs.jhu.edu
@@ 1 Introduction
While advanced LLMs perform well on standard MCQA benchmarks, such as MMLU Hendrycks et al. (2021), MathQA Amini et al. (2019), and GSM-8K Cobbe et al. (2021), they still face challenges when it comes to numerous choices that are closely related Wang et al. (2024a). In particular, we identify a critical problem: LLMs suffer from confidence distortion on MCQA tasks based on a comprehensive benchmark analysis. Specifically, LLM’s confidence in the correct choice tends to degrade, while its confidence in incorrect choices increases as the number of choices grows. This problem grows even more pronounced in the presence of additional noisy or partially relevant choices, thereby increasing the unreliable generations and erroneous predictions.
@@
To overcome this limitation, we introduce Self-Ensemble for LLM calibration in this work. Unlike existing work, Self-Ensemble calibrates LLM predictions without relying on a labeled validation dataset for parameter tuning. Specifically, Self-Ensemble mitigates the confidence distortion problem through a divide-and-conquer process: it splits the choices into several groups of choice subsets; then ensembles LLM predictions on each group to achieve the final decision. To plug into existing LLM architecture, we design attention masking and positional re-encoding mechanisms that enable the divide-and-conquer process to be executed internally during inference. Figure 1 shows the comprehensive performance of Self-Ensemble on three datasets and LLMs, where our method shows significant improvement over standard inference and baseline methods by addressing the confidence distortion problem. To summarize, the contributions of this work are as follows:
  • Confidence distortion. We identify the confidence distortion problem of LLMs in MCQA task, where LLMs have under-confident problems on correct answers and over-confident problems on incorrect answers.
@@ 2.2 LLMs for MCQA
3 Confidence Distortion and Self-Ensemble
We demonstrate the confidence distortion problem for LLMs and introduce our Self-Ensemble method to overcome this problem.
3.1 Confidence Distortion on Many-choice QA
We give Figure 2 to demonstrate the confidence distortion problems of LLMs on many-choice QA, where the experiment is on the QASC datasets Khot et al. (2020). Specifically, Figure 2 shows the proportion of $f([Q,\mathcal{C}])\!>\!\tau$ versus $\tau$ for $10\!<\!\tau\!<\!$ 1, where $f([Q,\mathcal{C}])\!>\!\tau$ represents LLM’s predicted probability on the correct or incorrect choice exceeds a threshold $\tau$. A high value of $f([Q,\mathcal{C}])\!>\!\tau$ proportion across $10\!\leq\!\tau\!\leq\!$ 1 indicates strong model confidence. To show the confidence distortion problem, we compare the $f([Q,\mathcal{C}])\!>\!\tau$ proportion across different choice settings with $K=2,4,6,8$ in Figure 2. On the correctly answered questions, the LLM achieves the highest $f([Q,\mathcal{C}])\!>\!\tau$ proportion when $K\!=\!2$, showing that the model becomes under-confident in many-choice scenarios. In contrast, for the incorrectly answered questions, the LLM shows the highest $f([Q,\mathcal{C}])\!>\!\tau$ proportion when $K\!=\!8$, indicating an over-confidence issue in many-choice scenarios.
The confidence distortion problem causes a degradation of LLM performance on many-choice QA. As shown in Figure 3, for the LLaMA-3-8B model, it can originally solve the 4-choice QA. However, when the same question is extended to 8 choices, while keeping the correct answer unchanged, the model selects an incorrect option. This failure is due to a decrease in confidence in the correct answer and an increase in confidence in incorrect options. We show a comprehensive result in Table 1, where each column shares the same question context but varies in the number of answer choices. A many-choice question contains all the options from the corresponding few-choice versions. For example, each question in the 8-choice task includes all the choices from the 6-choice task. As shown in Table 1, LLMs generally perform worse on many-choice QA than few-choice QA, This indicates the confidence distortion problem for LLM on many-choice QA.
@@ 3.2 LLM Inference with Self-Ensemble
We propose Self-Ensemble to solve the confidence distortion problem on many-choice QA. The intuition of Self-Ensemble is to divide a $K$-choice (many-choice) QA into multiple $m$-choice (few-choice) QA, where $m\!\ll\!K$. Then, it collects LLM’s answer probability in each few-choice case, and estimates the expected probability of each choice for the final decision. Specifically, Self-Ensemble has three steps as follows:
@@
$\!\!\!\!{Pr}(op_{i})\!=\!\mathbb{E}_{s\sim U(\mathbb{N}),{G}_{j}\sim\mathcal{G}_{s}(\mathcal{C}),\atop\widetilde{op}\sim{G}_{j}}\!\!\big{[}\mathbb{I}(op_{i}\!=\!\widetilde{op})Pr(\widetilde{op})\big{]}\!\!\!\!\!\vspace{-2mm}$ (2)
where $s\!\sim\!U(\mathbb{N}),{G}_{j}\!\sim\!\mathcal{G}_{s}(\mathcal{C})$ represents randomly splitting the choice set using different integer random seed $s$; and $\mathbb{I}(op_{i}\!=\!\widetilde{op})\!=\!1$ if $op_{i}\!=\!\widetilde{op}$; otherwise it takes 00. The intuition of Equation (2) is to estimate the expected probability of each choice by averaging over different group partitions.
@@ 4 Plugging Self-Ensemble into LLMs
We propose integrating Self-Ensemble into existing LLM architectures to enable intrinsic self-ensemble processing during inference. As illustrated in Equations (1) and (2), Self-Ensemble requires multiple forward passes through the LLM, which are $f([Q,{G}_{1}])$, $f([Q,{G}_{2}])$, $f([Q,{G}_{3}]),\cdots$. For efficient self-ensemble processing, we design specialized attention masks and positional encoding, such that these multiple calls of LLM can be executed in a single forward pass over the concatenated sequence $f([Q,{G}_{1},{G}_{2},{G}_{3},\cdots])$. Here, each group ${G}_{j}\sim\mathcal{G}_{s}(\mathcal{C}),s\sim\mathbb{N}$ is sampled from the choice set following Section 3.2 Step 1; and $[Q,{G}_{1},{G}_{2},{G}_{3},\cdots]$ denotes combining the question and different groups of choice into a sequence of tokens as prompts for LLMs.
@@ 4.1 Attention Masking
The standard auto-regressive attention mask follows $\mathbb{I}(j\leq i)$, ensuring that each token attends to all preceding tokens in the sequence. However, following Self-Ensemble, the LLM should restrict its attention to a single group of choices during each trial, ignoring the influence of other groups. Therefore, for the concatenated input sequence with the question and all choice groups $[Q,{G}_{1},{G}_{2},{G}_{3},\cdots]$, when processing a certain choice group $G_{i}$, the LLM should not have attention to other groups $G_{j}$, $j\neq i$.
@@
where $\texttt{Tk}(Q)$ denotes the tokens of question $Q$; $\mathbb{I}(\texttt{Tk}_{i},\!\texttt{Tk}_{j}\!\in\!\texttt{Tk}(G_{n}))\!=\!1$ if $\texttt{Tk}_{i}$ and $\texttt{Tk}_{j}$ belong to the same group, and 00 otherwise; and $M_{i,j}=0$ for $j>i$ as standard attention masks. This attention mask enforces pairwise independence between choice groups during inference.
@@ 4.2 Positional Re-encoding
The standard auto-regressive positional encoding follows $\texttt{Pos}_{i}\!=\!i$, where $\texttt{Pos}_{i}$ denotes the positional index of a token at position $i$. Following Self-Ensemble, the LLM should process each question-group pair $[Q,{G}_{j}]$ as a continuous input sequence for $1\!\leq\!j\!\leq\!N$. However, this continuity is disrupted in the concatenated input sequence $[Q,G_{1},G_{2},G_{3},\cdots]$, because starting from $G_{2}$, the question $Q$ and a choice group are no longer physically adjacent. For example, ${G}_{1}$ lies between $Q$ and ${G}_{2}$, breaking the positional continuity between $Q$ and $G_{2}$. To preserve the relative positional relationships between the question and each choice group, each choice group should be encoded into the relative contextual position to the question. Specifically, the positional indexes of question tokens and choice tokens are given as follows:
@@
Following Equation (4) to replace the position index of choice tokens, the position index of input sequence $[Q,{G}_{1},\cdots,{G}_{n}]$ is given by
@@
By combining the attention mask with positional re-encoding, we ensure that processing the concatenated sequence $[Q,G_{1},G_{2},G_{3},\cdots]$ is functionally equivalent to independently processing each question-group pair $[Q,G_{1}],[Q,G_{2}],[Q,G_{3}],\cdots$. This enables Self-Ensemble to seamlessly integrate into existing LLM architectures.
@@ 5 Experiment
In this experiment, we conduct experiments to evaluate Self-Ensemble, aiming to answer the following research questions: RQ1: Can Self-Ensemble improve on LLM’s accuracy on many-choice QA? RQ2: Can Self-Ensemble mitigate the confidence distortion problems for LLMs? RQ3: How does Self-Ensemble reshape the scaling of model accuracy with respect to parameter size?
@@ 5.2 Accuracy of MCQA (RQ1)
For datasets of different difficulty levels, standard inference shows less accuracy on TruthfulQA and MMLU-Pro than QASC. This implies that the confidence distortion is general problem on both easy and hard MCQA tasks. In contrast, Self-Ensemble consistently improves performance across all difficulty levels, demonstrating its general effectiveness in mitigating the confidence distortion problem.
@@ 5.3 Confidence Calibration by Self-Ensemble (RQ2)
In this section, we show that Self-Ensemble mitigates the confidence distortion problem of LLMs in Figure 6. Specifically, Figure 6 shows the proportion of $f([Q,\mathcal{C}])\!>\!\tau$ versus $\tau$ for $10\!<\!\tau\!<\!$ 1 on three different LLMs and datasets, where $f$ takes different LLMs; and $f([Q,\mathcal{C}])\!>\!\tau$ represents LLM’s predicted probability exceeds a threshold $\tau$. A higher value of $f([Q,\mathcal{C}])\!>\!\tau$ proportion across $10\!<\!\tau\!<\!$ 1 indicates strong model confidence.
@@ 5.5 Application to Quantized LLMs
To further explore Self-Ensemble’s application to memory-constrained settings, we show its effectiveness on quantized LLMs, including LLaMA-3-8B, Mistral-7B-v0.1, and Qwen2-7B under 4-bit bitsandbytes (BNB) quantization. As shown in Table 3, Self-Ensemble effectively offsets accuracy loss caused by the aggressive compression, enabling each quantized model to outperform its full-precision version. This indicates Self-Ensemble’s potential for quantized LLMs in resource-constrained environments, without requiring additional training or external information.

### v2

# Self-Ensemble: Mitigating Confidence Mis-calibration for Large Language Models

@@ Abstract
Although Large Language Models (LLMs) perform well in general fields, they exhibit a confidence mis-calibration problem on multi-choice question-answering (MCQA), particularly as the number of answer choices increases. Specifically, on MCQA with many choices, LLMs suffer from under-confidence in correct predictions and over-confidence in incorrect ones, leading to a substantially degraded performance. To solve this problem, we propose Self-Ensemble in this work. Our method splits the choices into several groups and ensembles LLM predictions across these groups to reach a final decision. The advantage of Self-Ensemble is its plug-and-play nature, where it can be integrated into existing LLM architecture based on a designed attention mask and positional encoding, without requiring labeled datasets for parameter tuning. Experimental results on three LLMs and datasets demonstrate that Self-Ensemble comprehensively addresses the confidence mis-calibration problem of LLMs, outperforming standard inference as well as baseline methods. The source code is available at https://github.com/ZichengXu/Self-Ensemble.
Self-Ensemble: Mitigating Confidence Mis-calibration for
@@
Zicheng Xu1,3thanks: Equal contribution, ordered by rolling dices., Guanchu Wang211footnotemark: 1, Guangyao Zheng3, Yu-Neng Chuang1, Alexander Szalay3, Xia Hu1, Vladimir Braverman3thanks: Correspondence to: Vladimir Braverman vova@cs.jhu .edu and Guanchu Wang gwang16@charlotte.edu 1Rice University, 2University of North Carolina at Charlotte, 3Johns Hopkins University
@@ 1 Introduction
While advanced LLMs perform well on standard MCQA benchmarks, such as MMLU Hendrycks et al. (2021), MathQA Amini et al. (2019), and GSM-8K Cobbe et al. (2021), they still face challenges when it comes to numerous choices that are closely related Wang et al. (2024a). In particular, we identify a critical problem: LLMs suffer from confidence mis-calibration on MCQA tasks based on a comprehensive benchmark analysis. Specifically, LLM’s confidence in the correct choice tends to degrade, while its confidence in incorrect choices increases as the number of choices grows. This problem grows even more pronounced in the presence of additional noisy or partially relevant choices, thereby increasing the unreliable generations and erroneous predictions.
@@
To overcome this limitation, we introduce Self-Ensemble for LLM calibration. Unlike existing work, Self-Ensemble calibrates LLM predictions without relying on a labeled validation dataset for parameter tuning. Specifically, Self-Ensemble mitigates the confidence mis-calibration through a divide-and-conquer process: it splits the choices into several groups of choice subsets; then ensembles LLM predictions on each group to achieve the final decision. To plug into existing LLM architecture, we design attention masking and positional re-encoding mechanisms that enable the divide-and-conquer process to be executed internally during inference. Figure 1 shows the comprehensive performance of Self-Ensemble on three datasets and LLMs, where our method shows significant improvement over standard inference and baseline methods by addressing the confidence mis-calibration. To summarize, the contributions of this work are as follows:
  • Confidence Mis-calibration. We identify the confidence mis-calibration problem of LLMs in MCQA task, where LLMs have under-confident problems on correct answers and over-confident problems on incorrect answers.
@@ 2.2 LLMs for MCQA
3 Confidence Mis-calibration and Self-Ensemble
We demonstrate the confidence mis-calibration for LLMs and introduce our Self-Ensemble method to overcome this problem.
3.1 Mis-calibration on Many-choice QA
We give Figure 2 to demonstrate the confidence mis-calibration of LLMs on many-choice QA, where the experiment is on the QASC datasets Khot et al. (2020). Specifically, Figure 2 shows the proportion of $f([Q,\mathcal{C}])\!>\!\tau$ versus $\tau$ for $0\!<\!\tau\!<\!1$, where $f([Q,\mathcal{C}])\!>\!\tau$ represents LLM’s predicted probability on the correct or incorrect choice exceeds a threshold $\tau$. A high value of $f([Q,\mathcal{C}])\!>\!\tau$ proportion across $0\!\leq\!\tau\!\leq\!1$ indicates strong model confidence. To show the confidence mis-calibration, we compare the $f([Q,\mathcal{C}])\!>\!\tau$ proportion across different choice settings with $K=2,4,6,8$ in Figure 2. On the correctly answered questions, the LLM achieves the highest $f([Q,\mathcal{C}])\!>\!\tau$ proportion when K=2K\!=\!2, showing that the model becomes under-confident in many-choice scenarios. In contrast, for the incorrectly answered questions, the LLM shows the highest $f([Q,\mathcal{C}])\!>\!\tau$ proportion when K=8K\!=\!8, indicating an over-confidence issue in many-choice scenarios.
The problem of confidence mis-calibration causes a degradation of LLM performance on many-choice QA. As shown in Figure 3, for the LLaMA-3-8B model, it can originally solve the 4-choice QA. However, when the same question is extended to 8 choices, while keeping the correct answer unchanged, the model selects an incorrect option. This failure is due to a decrease in confidence in the correct answer and an increase in confidence in incorrect options. We show a comprehensive result in Table 1, where each column shares the same question context but varies in the number of answer choices. A many-choice question contains all the options from the corresponding few-choice versions. For example, each question in the 8-choice task includes all the choices from the 6-choice task. As shown in Table 1, LLMs generally perform worse on many-choice QA than few-choice QA, This indicates the confidence mis-calibration for LLM on many-choice QA.
@@ 3.2 LLM Inference with Self-Ensemble
We propose Self-Ensemble to address the confidence mis-calibration on many-choice QA. The intuition of Self-Ensemble is to divide a $K$-choice (many-choice) QA into multiple $m$-choice (few-choice) QA, where $m\!\ll\!K$. Then, it collects LLM’s answer probability in each few-choice case, and estimates the expected probability of each choice for the final decision. Specifically, Self-Ensemble has three steps as follows:
@@
[…]
where $s\!\sim\!U(\mathbb{N}),{G}_{j}\!\sim\!\mathcal{G}_{s}(\mathcal{C})$ represents randomly splitting the choice set using different integer random seed $s$; and $\mathbb{I}(op_{i}\!=\!\widetilde{op})\!=\!1$ if $op_{i}\!=\!\widetilde{op}$; otherwise it takes 0. The intuition of Equation (2) is to estimate the expected probability of each choice by averaging over different group partitions.
@@ 4 Plugging Self-Ensemble into LLMs
We propose integrating Self-Ensemble into existing LLM architectures to enable intrinsic self-ensemble processing during inference. As illustrated in Equations (1) and (2), Self-Ensemble requires multiple forward passes through the LLM, which are $f([Q,{G}_{1}])$, $f([Q,{G}_{2}])$, $f([Q,{G}_{3}]),\cdots$. For efficient self-ensemble processing, we design specialized attention masks and positional encoding, such that these multiple calls of LLM can be executed in a single forward pass over the concatenated sequence $f([Q,{G}_{1},{G}_{2},{G}_{3},\cdots])$. Here, each group ${G}_{j}\sim\mathcal{G}_{s}(\mathcal{C}),s\sim\mathbb{N}$ is sampled from the choice set following Section 3.2 Step 1; and $Q,{G}_{1},{G}_{2},{G}_{3},\cdots]$ denotes combining the question and different groups of choice into a sequence of tokens as prompts for LLMs.
@@ 4.1 Attention Masking
The standard auto-regressive attention mask follows $\mathbb{I}(j\leq i)$, ensuring that each token attends to all preceding tokens in the sequence. However, following Self-Ensemble, the LLM should restrict its attention to a single group of choices during each trial, ignoring the influence of other groups. Therefore, for the concatenated input sequence with the question and all choice groups $Q,{G}_{1},{G}_{2},{G}_{3},\cdots]$, when processing a certain choice group $G_{i}$, the LLM should not have attention to other groups $G_{j}$, $j\neq i$.
@@
where $\texttt{Tk}(Q)$ denotes the tokens of question $Q$; $\mathbb{I}(\texttt{Tk}_{i},\!\texttt{Tk}_{j}\!\in\!\texttt{Tk}(G_{n}))\!=\!1$ if $\texttt{Tk}_{i}$ and $\texttt{Tk}_{j}$ belong to the same group, and 0 otherwise; and $M_{i,j}=0$ for $j>i$ as standard attention masks. This attention mask enforces pairwise independence between choice groups during inference.
@@ 4.2 Positional Re-encoding
The standard auto-regressive positional encoding follows $\texttt{Pos}_{i}\!=\!i$, where $\texttt{Pos}_{i}$ denotes the positional index of a token at position $i$. Following Self-Ensemble, the LLM should process each question-group pair $Q,{G}_{j}]$ as a continuous input sequence for $1\!\leq\!j\!\leq\!N$. However, this continuity is disrupted in the concatenated input sequence $Q,G_{1},G_{2},G_{3},\cdots]$, because starting from $G_{2}$, the question $Q$ and a choice group are no longer physically adjacent. For example, ${G}_{1}$ lies between $Q$ and ${G}_{2}$, breaking the positional continuity between $Q$ and $G_{2}$. To preserve the relative positional relationships between the question and each choice group, each choice group should be encoded into the relative contextual position to the question. Specifically, the positional indexes of question tokens and choice tokens are given as follows:
@@
Following Equation (4) to replace the position index of choice tokens, the position index of input sequence $Q,{G}_{1},\cdots,{G}_{n}]$ is given by
@@
By combining the attention mask with positional re-encoding, we ensure that processing the concatenated sequence $Q,G_{1},G_{2},G_{3},\cdots]$ is functionally equivalent to independently processing each question-group pair $Q,G_{1}],[Q,G_{2}],[Q,G_{3}],\cdots$. This enables Self-Ensemble to seamlessly integrate into existing LLM architectures.
@@ 5 Experiment
In this experiment, we conduct experiments to evaluate Self-Ensemble, aiming to answer the following research questions: RQ1: Can Self-Ensemble improve on LLM’s accuracy on many-choice QA? RQ2: Can Self-Ensemble mitigate the confidence mis-calibration problems for LLMs? RQ3: How does Self-Ensemble reshape the scaling of model accuracy with respect to parameter size?
@@ 5.2 Accuracy of MCQA (RQ1)
For datasets of different difficulty levels, standard inference shows less accuracy on TruthfulQA and MMLU-Pro than QASC. This implies that the confidence mis-calibration is a general problem on both easy and hard MCQA tasks. In contrast, Self-Ensemble consistently improves performance across all difficulty levels, demonstrating its general effectiveness in mitigating the confidence mis-calibration problem.
@@ 5.3 Confidence Calibration by Self-Ensemble (RQ2)
In this section, we show that Self-Ensemble mitigates the confidence mis-calibration of LLMs in Figure 6. Specifically, Figure 6 shows the proportion of $f([Q,\mathcal{C}])\!>\!\tau$ versus $\tau$ for $0\!<\!\tau\!<\!1$ on three different LLMs and datasets, where $f$ takes different LLMs; and $f([Q,\mathcal{C}])\!>\!\tau$ represents LLM’s predicted probability exceeds a threshold $\tau$. A higher value of $f([Q,\mathcal{C}])\!>\!\tau$ proportion across $0\!<\!\tau\!<\!1$ indicates strong model confidence.
@@ 5.4 Improving Accuracy-to-Parameter Scaling (RQ3)
5.5 General-purpose LLMs with Self-Ensemble Outperforms Reasoning LLMs and Chain of Thought Prompting
We evaluate whether Self-Ensemble can match or surpass specialized reasoning methods. On the QASC dataset, general-purpose LLMs combined with Self-Ensemble outperform their reasoning-tuned counterparts, indicating that Self-Ensemble unlocks accuracy gains without dedicated reasoning fine-tuning. Furthermore, Self-Ensemble also surpasses chain-of-thought (CoT) prompting Wei et al. (2022), which yields higher accuracy than CoT while avoiding the extra latency and token-generation overhead of producing reasoning traces. Both results are summarized in Table 3, highlighting that Self-Ensemble offers an efficient, single-pass, and fine-tuning-free alternative to reasoning LLMs and CoT prompting on MCQA.
Model Accuracy
LLaMA-3-8B 73.22
DeepSeek-R1-Distill-Llama-8B 73.33
LLaMA-3-8B with CoT 74.62
LLaMA-3-8B with Self-Ensemble 80.67
Qwen2-7B 69.44
DeepSeek-R1-Distill-Qwen-7B 70.63
Qwen2-7B with CoT 74.19
Qwen2-7B with Self-Ensemble 77.86
@@ 5.6 Application to Quantized LLMs
To further explore Self-Ensemble’s application to memory‐constrained settings, we show its effectiveness on quantized LLMs, including LLaMA-3-8B, Mistral-7B-v0.1, and Qwen2-7B under 4-bit bitsandbytes (BNB) quantization. As shown in Table 4, Self-Ensemble effectively offsets accuracy loss caused by the aggressive compression, enabling each quantized model to outperform its full-precision version. This indicates Self-Ensemble’s potential for quantized LLMs in resource-constrained environments, without requiring additional training or external information.
@@ 5.7 Ablation Study
5.8 Sensitivity to Group Size and Number of Trials
We study how Self-Ensemble depends on the group size $m$ and the number of ensembles $N$. Using Qwen2.5-7B on the QASC dataset, we sweep $m\in\{3,4,5\}$ and $N\in\{10,20,40,80\}$. As shown in Table 6, accuracy is relatively insensitive to the group size, while increasing the number of trials improves performance and stabilizes at $N\geq 40$, indicating robustness to hyperparameters. Consequently, Self-Ensemble can be used without hyperparameter tuning, reducing setup time while preserving performance.
Trials NN $m=3$ $m=4$ $m=5$
10 78.73 78.08 78.61
20 79.59 80.67 79.91
40 79.27 81.97 81.43
80 80.45 82.40 81.97
Table 6: Accuracy comparison of Self-Ensemble across group sizes and trial numbers.